May 17, 2018
Meeting URL https://bluejeans.com/388466836
Meeting ID 388 466 836
US toll free phone 1.888.240.2560
Dial in, type in the meeting ID and then press ##
1. GEANT4 simulation update -Eric K.[ slides]
2. GEANT3 simulation update -Bill H.[ slides]
3. Discussion of Kerr activities
Discussion of GEANT4. No new results to show. Eric is working on understanding seemingly unrealistically small weights on generated events.
Discussion of GEANT3. No new results to show. Working on generating more optimal quad settings for various beam energies.
Miscellany Dave Gaskell has not heard from D. Meekins on the target ladder but will inquire this week. Sasha reports that he is making progress on TOSCA model and will soon have a simple model for the entire set of quads and dipole. He expects to have results to show in two weeks! Roman will be leaving the country in two days for an undetermined length of time. Simona Malace will be taking over his studies of the new FADC-based DAQ.
Kerr discussion. Jinlong will be traveling to the lab the first week of June. If the magnet is ready for use he will prolong his stay for Kerr studies. Otherwise, if the magnet is not expected to be in service for a few weeks, he will transfer the equipment back to Stony Brook and continue studies there.
Email from Javier on magnet status "The magnet has been moved from the Test Lab to the Experiment Staging Building. There, we tried several times but have not been able to find any external leak on the magnet that could explain the condensation observed on the target shaft while warming the magnet on the Test Lab. This leaves us with at least two other possibilities - the leak is internal (in the pulse tube refrigerator) as suggested by Chris or we have not been reaching the proper magnet vacuum level before starting to cool down.
The magnet continues to be pumped down. The Hall A techs need the pumping station/leak checker which is connected to the magnet now so we are looking to borrow/replace it with another pumping station today/tomorrow. We will continue pumping until ready to start cooling down the magnet - we need a pressure of at least 10-4 Torr to begin. Electrical work is still needed to be able to power the compressor/water chiller - hopefully the work will get done by next week. The accelerator network installation to monitor the magnet while it cools down and to test the target motion mechanics should also be installed by then "